OpenStudy (twizttiez):

What is the best approximation of the solution to the system to the nearest integer values? A. (1, 3) B. (3, 2) C. (2, 3) D. (3, 4)

OpenStudy (twizttiez):

@princeharryyy

OpenStudy (twizttiez):

Soooo where does that leave me xD

OpenStudy (princeharryyy):

(2, 3)

OpenStudy (princeharryyy):

@TwiztTiez

OpenStudy (twizttiez):

Do you know what to do @princeharryyy

OpenStudy (twizttiez):

1 more

OpenStudy (princeharryyy):

recon solved it for u x=24/13 ~2 and y =132/39 ~3

OpenStudy (princeharryyy):

~ means nearly equal to ...

OpenStudy (princeharryyy):

u jst need to solve those 2 given eqns

OpenStudy (princeharryyy):

after solving those 2 equations... x=24/13 nearly equal to 2 and y =132/39 nearly equal to 3

OpenStudy (princeharryyy):

do u understand @TwiztTiez

OpenStudy (princeharryyy):

yup! quickly eyes r getting heavier..)))

OpenStudy (princeharryyy):

@TwiztTiez

OpenStudy (twizttiez):

Ok thanks one second!